Circular No. 8603

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005ed, 2005ee, 2005ef, 2005eg, 2005eh, 2005ei Six supernovae have been reported: 2005ed and 2005ef–2005ei via J.

Frieman on behalf of the “Sloan Digital Sky Survey II” collaboration (J. Barentine et al.; astrometric and photometric data appear on CBET 229), on g′, r′, and i′ images taken on Sept. 9 with the SDSS 2.5-m telescope at Apache Point Observatory; and 2005ee via V. Lipunov on behalf of his group at the Sternberg Astronomical Institute (cf. IAUC 8520) on unfiltered CCD MASTER telescope images. Lipunov writes that 2005ee is located at α = 23^{h}57^{m}55^{s}.83, δ = +32^{o}38′08′′.9 (equinox 2000.0), or 3′′ west and 5′′ north of the center of PGC 73054. Unfiltered CCD magnitudes for 2005ee: Aug. 25.92 UT, 16.0 (Lipunov et al.); Sept. 1.921, 16.0 (Lipunov et al.); 23.26, 16.6 (J. McGaha, Tucson, AZ, 0.36-m reflector); 24.02, 16.8 (T. Puckett, remotely with 0.20-m reflector at Osyoss, BC). Alternate measured position end figures for 2005ee: 55^{s}.97, 14′′.1 (McGaha); 55^{s}.88, 14′′.5 (Puckett). McGaha and Puckett each indicate that 2005ee is not present on Digitized Sky Survey images. Spectroscopy for the SDSS supernovae (each of which was predicted to have peaked in the g′-magnitude range 18.9–20.5 prior to discovery) show four of them to be of type Ia and the fifth (2005ei) a likely type-Ia supernova.

SUPERNOVA 2005cs IN NGC 5194 (M51) C. J. Stockdale and M. Kelley, Marquette University; S. D. Van Dyk,

Spitzer Science Center, California Institute of Technology; R. A. Sramek, National Radio Astronomy Observatory; K. W. Weiler, Naval Research Laboratory; and N. Panagia, European Space Agency, report the lack of detection of radio emission near the position of the type-II SN 2005cs (IAUC 8553, 8555) with the Very Large Array. Upper limits (3σ) to any radio flux density were established on the following dates: July 2.01 UT, < 0.585 mJy at 22.460 GHz (wavelength λ = 1.3 cm); 3.00, < 0.900 mJy at 14.940 GHz (λ = 2.0 cm) and < 0.189 mJy at 8.460 GHz (λ = 3.5 cm); 8.968, < 0.291 mJy at 8.460 GHz and < 0.918 mJy at 22.460 GHz; 22.038 < 0.161 mJy at 8.460 GHz and < 0.464 mJy at 22.460 GHz; Aug. 9.00, < 0.129 mJy at 8.460 GHz; 14.837, < 0.187 mJy at 8.460 GHz and < 0.372 mJy at 4.860 GHz (λ = 6.2 cm). The search for radio emission was conducted within ∼ 10′′ of the published optical position of SN 2005cs (α = 13^{h}29^{m}52^{s}.78, δ = +47^{o}10′35′′.7, equinox 2000.0; cf. IAUC 8553), and no radio emission was detected near the supernova position exceeding three times the map rms at any of the times or in any of the frequency bands listed.

2005 September 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green